OpenStudy (calculusxy):

If the average of four consecutive odd numbers is 24, what is the largest number? (A) 23 (B) 25 (C) 27 (D) 29 (E) 96

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

do you have the equation for this? or are you stuck on that part?

OpenStudy (calculusxy):

\[\frac{ x + (x+2) + (x+4) + (x+6) }{ 4 } = 96\]

OpenStudy (calculusxy):

sorry 24

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

oh right, yes, the right side is 24

OpenStudy (solomonzelman):

well 96 is after you ×4 on both sides;)

OpenStudy (calculusxy):

i mean i multiple 4 times 24, which equals 96

OpenStudy (calculusxy):

@SolomonZelman yes exactly

OpenStudy (solomonzelman):

then this should not be a prob for you. good luck:)

OpenStudy (calculusxy):

\[4x + 12 = 96 => 4x =84 => x = 21\]

OpenStudy (calculusxy):

and then 21 + 6 = 27

jimthompson5910 (jim_thompson5910):

x = 21 is correct and yes, the largest number is 27

OpenStudy (calculusxy):

thank you again!

jimthompson5910 (jim_thompson5910):

sure thing